A car is traveling at 55.0 mi/h on a horizontal highway.

(a) If the coefficient of static friction between road and tires on a rainy day is 0.102, what is the minimum distance in which the car will stop?


(b) What is the stopping distance when the surface is dry and ?s = 0.605?
Answer #1

here,

intial speed of car , u = 55 mi/h = 24.6 m/s

a)

us = 0.102

accelration , a = - us * g

a = - 0.102 * 9.81 = - 1 m/s^2

let the minimum stopping distance be s

v^2 - u^2 = 2 * a * s

0 - 24.6^2 = - 2 * 1 * s

s = 302.6 m

b)

us = 0.0.605

accelration , a = - us * g

a = - 0.605 * 9.81 = - 6.38 m/s^2

let the minimum stopping distance be s

v^2 - u^2 = 2 * a * s

0 - 24.6^2 = - 2 * 6.38 * s

s = 47.5 m


answered by: ANURANJAN SARSAM